OpenStudy (anonymous):

What is the equation of the line meeting the given conditions.(In standard form) Containing C(3, 3) and D(-3, 2)

OpenStudy (anonymous):

First find the slope, do you know how to do that?

OpenStudy (anonymous):

Between these two points? Yes.

OpenStudy (anonymous):

Ok compute the slope, then use the point slope equation: y-y1=m(x-x1), where m is the slope and (x1,y1) is a given coordinate

OpenStudy (anonymous):

Wait which y is first y2?

OpenStudy (anonymous):

for the slope equation? or the point slope equation?

OpenStudy (anonymous):

point slope

OpenStudy (anonymous):

2-3=1/6(-3-3) is that it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Forget it I found a calculator. Lol. Thanks for you time though!